OpenStudy (anonymous):

f(t)=Q0(1+r)^t. Find the growth rate, r, to the nearest thousandth, given f(0.01)=1.06 and f(0.11)=1.09.

OpenStudy (campbell_st):

well substituting you have \[1.06 = Q_{0}( 1 + r)^{0.01}\] and you also have \[1.09 = Q_{0}(1 + r)^{0.11}\] using a little index law property about adding powers when multiplying you can rewrite the 2nd equation as \[1.09 = Q(1 + r)^{0.01} \times (1 + r)^{0.1}\] now you can do a little substitution from the 1st equation \[1.09 = 1.06 \times (1 + r)^{0.1}\] divide both sides of the equation by 1.06 \[\frac{1.09}{1.06} = (1 + r)^{0.1}\] since 0.1 = 1/10 you have an equation where \[\frac{1.09}{1.06} = ( 1 + r)^{\frac{1}{10}}\] raise each term to the 10th power, then \[(\frac{1.09}{1.06})^{10} = 1 + r\] this should make it easy to find the growth rate r...